OpenStudy (anonymous):

Please help me create an explicit and recursive formula for the geometric sequence -6, 12, -24, ... with a common ratio of -2 so that I can find a sub 10!

OpenStudy (anonymous):

\[-6,-6\times (-2), -6\times (-2)^2,-6\times (-2)^3,...\] general formula will be \[a_n=-6\times (-2)^{n-1}\]